In 2014, a farmer grows and sells $3 million worth of corn to Big Flakes Cereal Company. Big Flakes Cereal Company produces $8 million worth of cereal in 2014, with sales to households during the year of $7 million. The unsold $1 million worth of cereal remains in Big Flake Cereal Company’s inventory at the end of 2014. The transactions just described contribute how much to GDP for 2014?
Answer #1

Value added by the farmer = 3 million

Value added by the big cereal company = 8-3 = 5 million

Total value added to the GDP in 2014 = 3+5 = $8 million

GDP is the value of total output produced during the given year